OpenStudy (anonymous):

ANSWER THESE QUESTIONS!!!!!! :D ANSWER THESE QUESTIONS!!!!!! :D @Mathematics

OpenStudy (anonymous):

Is this your homework?

OpenStudy (anonymous):

Is there a reason you need them solved?

OpenStudy (anonymous):

Then if it's not your homework, what's the reason? ;)

OpenStudy (anonymous):

Which ones are you specifically struggling with?

OpenStudy (anonymous):

2b and 2c 3a 4, 8

OpenStudy (anonymous):

(b) 3t 2 + t = 4 7t. (c) 3y5 6 = 2y1 3 + 2

OpenStudy (anonymous):

Problem 2 (b and c) Problem 3 (a)Problem 4 (all) and Problem 8 (all)

OpenStudy (anonymous):

2b: \[\Large \begin{array}{l} 3t - 2 + t = 4 - 7t\\ 3t + t + 7t = 4 + 2\\ 11t = 6\\ t = \frac{6}{{11}} \end{array}\]